# Private Counterfactual Retrieval

### v1

@@ 3.3 Mask-PCR
There exists a scheme that has a lower leakage in terms of (5) compared to the baseline PCR with communication cost $2(d+M)$.
@@
$d_{k}=\min_{i,j}|d_{i}(x_{k})-d_{j}(x_{k})|,$ (21)
$d_{\min}=\min_{k}d_{k}.$ (22)
Let $\mu$ be a random variable with support $\{0,\ldots,d_{\min}-1\}$. Now, a user who wishes to know the closest accepted sample to their rejected sample $x\in\mathcal{D}_{c}$ sends the query in (12) to the $n$th server.
Upon receiving the queries, each server computes the answers as follows
$\!\!A_{n}^{[x]}(i)=||y_{i}-Q_{n}^{[x]}||^{2}+\mu(i)+\alpha_{n}Z^{\prime}(i),i\in[M],\!$ (23)
where $\mu(i)$ has the same distribution as $\mu$ and $Z^{\prime}(i)$ is a uniform random variable in $\mathbb{F}_{q}$.
As the user receives the answers from the servers, they are reprocessed as follows
@@
then use them to decode the distances as follows
@@
Upon getting the values $d_{i}(x)+\mu(i)$, $i\in[M]$, the user utilizes them to decide which is closest based on their numerical value, i.e., decode the index of the closest accepted sample. To show that the user can decode correctly although the exact distances are masked, the following lemma is provided.
@@
Note that the user does not know the exact value of $d_{\min}$ since they do not have any prior knowledge of the accepted samples.
Communication Cost:
@@
Field size Effect in Masking
The field size can have a significant role in counterfactual retrieval with the masking approach. Recall that the minimum requirement for the field size is $R^{2}d$ as explained at the beginning of this section. Let $q_{1}$ be the field size satisfying $q_{1}>R^{2}d$. If another field size $q_{2}$ is chosen such that $q_{2}>q_{1}$, we can embed the samples $y_{1},\ldots,y_{M}$ in $\mathbb{F}_{q_{2}}$ using a relative distance preserving transform $T:\mathbb{F}_{q_{1}}\rightarrow\mathbb{F}_{q_{2}}$ such that the following condition is satisfied
$|d_{i}(x_{k})-d_{j}(x_{k})|\leq$ $|\big{(}||x_{k}-T(y_{i})||^{2}-||x_{k}-T(y_{j})||^{2}\big{)}|,$
$\forall k,i,j.$ (26)
@@
This example demonstrates the masking approach and the field size expansion. Let the rejected samples, for example, be $\{[1,2]^{t},[2,1]^{t}\}$, and the set of accepted samples be $\{[20,0]^{t},[0,20]^{t}\}$. Let the field size be $q_{1}=401$. Thus, after simple calculations, we see that the range of the masking random variable is $\{0,\ldots,39\}$. Let the user choose $x=[1,2]^{t}$. Thus, the queries sent to the two servers are given by (12) and after the servers reply and the user reprocesses the received samples as mentioned previously, the answers are as follows,
$A^{[x]}(1)=\begin{bmatrix}A_{1}^{[x]}(1)\\ A_{2}^{[x]}(1)\end{bmatrix}=\bm{M}_{2}\begin{bmatrix}365+\mu(1)\\ I(1)+Z^{\prime}(1)\end{bmatrix},$ (27)
@@
$A^{[x]}(2)=\begin{bmatrix}A_{1}^{[x]}(2)\\ A_{2}^{[x]}(2)\end{bmatrix}=\bm{M}_{2}\begin{bmatrix}325+\mu(2)\\ I(2)+Z^{\prime}(2)\end{bmatrix}.$ (28)
It is clear that for any choice of $\mu(1)$, and $\mu(2)$, the user would be able to know that $y_{1}$ is closest to $x$.
Now, let us choose a larger field size, for example, $q_{2}=40009$ and apply the simple transform $T:a\rightarrow 10a$. Then, with simple calculations, we see that the range of the masking random variable is $\{0,399\}$ which is larger than the previous one and still maintains the relative distance.
Leakage Analysis:
The leakage analysis is presented in Appendix B.
4 PCR WITH ACTIONABILITY
In this section, we assume that the user assigns different weights to their attributes based on their preference to change those attributes to attain a counterfactual. The more reluctant the user is to change a given attribute, the higher its weight is. In order to ensure that the user’s preference is also private, the weight vector should be kept private from the servers. We present achievable schemes that require $N=3$ replicated databases.
4.1 Baseline PCR+
@@
There exists a scheme that can retrieve the exact closest counterfactual with user’s actionability weights using $N=3$ servers while keeping the user’s sample and actionability weights hidden from each server. In addition, the total communication cost of this scheme is $3(2d+M)$.
To provide the achievability scheme for Theorem 5, let the weight vector be denoted as $w\in[L]^{d}$. In this case, the required minimum field size is $q\geq R^{2}Ld$. Now, the user sends the query tuple $Q_{n}^{[x,w]}$ to server $n$, where
$Q_{n}^{[x,w]}(1)=x+\alpha_{n}Z_{1},$ (29)
$Q_{n}^{[x,w]}(2)=w+\alpha_{n}Z_{2},$ (30)
@@
$-2(y_{i}-x)^{t}\bm{W}Z_{1}+Z^{\prime}_{1}\Big{)}-\alpha_{n}^{2}\Big{(}Z_{1}^{t}\bm{W}Z_{1}$
$+2(y_{i}-x)^{t}\bm{Z}Z_{1}+Z^{\prime}_{2}\Big{)}+\alpha_{n}^{3}Z_{1}^{t}\bm{Z}Z_{1},$ (32)
@@
Upon receiving the answers from $N=3$ servers, the user applies the following decoding approach
@@
$=\bm{M}_{3}\begin{bmatrix}d_{w}(y_{i},x)\\ I_{1}(i)\\ I_{2}(i)\end{bmatrix},$ (34)
@@
4.2 Diff-PCR+
Theorem 6
There exists a scheme that provides lower leakage compared to the baseline PCR+ scheme with a communication cost $3(2d+M-1)$.
@@ 4.2 Diff-PCR+
where $I_{1}^{\prime}(i)=(y_{i}-x)^{t}\bm{Z}(y_{i}-x)-(y_{i+1}-x)^{t}\bm{Z}(y_{i+1}-x)+2Z_{1}^{t}\bm{W}(y_{i}-y_{i+1})+Z^{\prime}_{1}(i)$ and $I^{\prime}_{2}(i)=2Z_{1}^{t}\bm{Z}(y_{i}-y_{i+1})+Z^{\prime}_{2}(i)$ are the interference terms. The answers $A_{n}^{[x,w]}(i)$ of the three servers can be written as,
$\begin{bmatrix}A_{1}^{[x,w]}(i)\\ A_{2}^{[x,w]}(i)\\ A_{3}^{[x,w]}(i)\end{bmatrix}=\bm{M}_{3}\begin{bmatrix}d_{w}(y_{i},x)-d_{w}(y_{i+1},x)\\ I_{1}^{\prime}(i)\\ I_{2}^{\prime}(i)\end{bmatrix}.$ (39)
Therefore, the user recovers the $M-1$ differences, while discarding the interference terms.
Finding the Closest $y_{i}$ to $x$:
Now, with the $M-1$ differences, the user evaluates their counterfactual index $\theta^{*}$ following sequential comparisons similar to Algorithm 1, the only difference being the range in line 5 is replaced by $\mathcal{H}^{c}$.
Communication Cost:
@@
4.3 Mask-PCR+
Theorem 7
There exists a scheme that satisfies the optimization problem defined in (6) with weighted preferences being private from the servers. In addition, it provides less leakage compared to the baseline PCR+ scheme. The communication cost of this scheme is $3(2d+M)$.
The queries are the same as in the previous sections. Upon receiving the queries, each server applies the following on each sample $y_{i}$ and sends to the user
@@ 4.3 Mask-PCR+
where $\mu(i)$ is as defined in the previous section, $Z^{\prime}_{1}(i)$, and $Z^{\prime}_{2}(i)$ are uniform random variables and $I_{1}(i)$ and $I_{2}(i)$ are as given in (4.1). Using the answers of the $N=3$ servers and applying the decoding approach given in Section 4.1, the user obtains the masked weighted distance corresponding to $y_{i}$ as follows:
$\hat{A}^{[x,w]}(i)=\bm{M}_{3}\begin{bmatrix}d_{w}(y_{i},x)+\mu(i)\\ I_{1}(i)\\ I_{2}(i)\end{bmatrix}.$ (42)
To show that there is no need to change the range of the masking random variable $\mu$ with any weighting, we provide the following lemma.
@@
Proof:  Note that,
$|(y_{i}-x_{k})^{t}\bm{W}(y_{i}-x_{k})-(y_{j}-x_{k})^{t}\bm{W}(y_{j}-x_{k})|$
$=|\sum_{\ell}w_{\ell}(y_{i}(\ell)-x_{k}(\ell))^{2}-w_{\ell}(y_{j}(\ell)-x_{k}(\ell))^{2}|$ (43)
$\geq|\sum_{\ell}(y_{i}(\ell)-x_{k}(\ell))^{2}-(y_{j}(\ell)-x_{k}(\ell))^{2}|$ (44)
$=|d_{w}(y_{i},x_{k})-d_{w}(y_{j},x_{k})|.$ (45)
Thus, $\min_{k}\min_{i,j}|(y_{i}-x_{k})^{t}\bm{W}(y_{i}-x_{k})-(y_{j}-x_{k})^{t}\bm{W}(y_{j}-x_{k})|\geq d_{\min}$. $\blacksquare$
Communication Cost:
@@
5 EXPERIMENTS
5.1 Accuracy-Quantization Trade-Off
As detailed in Section 3.1, the proposed schemes operate on finite fields with size $q$. However, a real-world application might require some of the features to be real-valued. To circumvent this technicality, we may quantize each feature of the counterfactual instances as well as the user queries before applying the schemes. However, the quantization will lead to a loss of some information that might result in an error in finding the closest counterfactual. The error can be made arbitrarily small by increasing the number of quantization levels, as we demonstrate in the following experiments. The trade-off is that a larger finite field $\mathbb{F}_{q}$ will be required for the scheme to work properly; e.g., for two implementations of Mask-PCR with numbers of quantization levels $R_{1},R_{2}$ such that $R_{1}>R_{2}$, the minimum field sizes required are $q_{1}\geq q_{2}$ because $q_{1}>R_{1}^{2}d$ and $q_{2}>R^{2}_{2}d$.
Moreover, in the case of Mask-PCR, we may lift the restriction $x\in\mathcal{D}_{c}$ off by empirically deciding the parameter $d_{\min}$ to be used such that the scheme works for most of the points in a sample set of queries. As pointed out in Section 3.3, a larger field size $q$ is needed to accommodate a larger value for $d_{\min}$.
We carry out our experiment using the Wine Quality Dataset (Cortez et al., 2009) to observe the accuracy-quantization trade-offs mentioned above. The dataset contains 4898 instances, each with 11 real-valued features. The original target variable “Quality” is categorical, taking integer values from 0 to 10. We convert the target into a binary variable by defining $T=\mathds{1}[``Quality"\geq 5]$. This gives us 3788 instances with $T=1$ and 183 instances with $T=0$ (which are the potential queries).
We consider all the instances with $T=1$ to be the accepted instances (potential counterfactual instances). Let this set be $\mathcal{S}_{1}$. The set $\mathcal{S}_{0}$ consists of all the instances with $T=0$ which are the potential queries. In each round, we pick a subset $\mathcal{D},(|\mathcal{D}|=M)$ of $\mathcal{S}_{1}$ as the database and a subset $\mathcal{S}_{x}$ of $\mathcal{S}_{0}$ as the queries. Then, we apply the retrieval schemes with varying parameters to obtain the nearest neighbor counterfactuals. The accuracy of the retrieved counterfactuals is computed as follows: Denote the counterfactual of the query $x$ retrieved using the scheme by $\hat{y}(x)$ and the actual (i.e., the exact nearest neighbor) counterfactual by $y(x)$. Note that the scheme provides the index of the counterfactual, hence, we can retrieve the exact unquantized counterfactual irrespective of the quantization applied. The accuracy of the scheme is
@@ 5.1 Accuracy-Quantization Trade-Off
Results of the experiment, in Figure 2, show how the accuracy improves with an increasing number of quantization levels. Further, for a given number of quantization levels, the accuracy degrades as $d_{\min}$ increases.
@@
5.2 Database Leakage Results
To observe the success of Diff-PCR in mitigating the leakage, we compute the exact leakage values over a synthetic dataset. We consider $R=4$ and $d=2$ with a database size of $M=5$. We assume the queries are equi-probable over the $[0:R]^{d}$ space. We further assume that given the query $x$, $y_{1},\dots,y_{M}$ are equi-probable over $[0:R]^{d}\backslash\{x\}$. Table 1 lists the computed leakage values under these assumptions.
Scheme Leakage (to base $q$)
Baseline PCR 3.1297
Diff-PCR 2.0404
Table 1: Leakage results.
Appendix A Proof of Lemma 1
Proof:  First, assume that $x\in\{x_{1},\ldots,x_{K}\}$. To prove the required result, it suffices to prove that the random variable used preserves the order of the distances between samples, i.e., $d_{i}(x)-d_{j}(x)>0$ if $y_{j}$ is closer to $x$ and $d_{i}(x)-d_{j}(x)<0$ if $y_{i}$ is closer to $x$. Let $|d_{i}(x)-d_{j}(x)|=\ell$. Then,
$d_{i}(x)-d_{j}(x)=\begin{cases}\ell,&\text{$y_{j}$ is closer to $x$},\\ -\ell,&\text{$y_{i}$ is closer to $x$}.\end{cases}$ (47)
Now, assume $y_{j}$ is closer, thus
$d_{i}(x)-d_{j}(x)+\mu(i)-\mu(j)$ $=\ell+\mu(i)-\mu(j)$
$\geq\ell-(d_{\min}-1)$
$>0,$ (48)
and if $y_{i}$ is closer, we have
$d_{i}(x)-d_{j}(x)+\mu(i)-\mu(j)$ $=-\ell+\mu(i)-\mu(j)$
$\leq-\ell+(d_{\min}-1)$
$<0.$ (49)
Now, assume $x\in\mathcal{D}_{c}$. Using the same approach as above, and the definition of $x\in\mathcal{D}_{c}\setminus\{x_{1},\ldots,x_{K}\}$, we have $|d_{i}(x)-d_{j}(x)|\geq d_{\min}$, which proves the required result. $\blacksquare$
Appendix B Mask-PCR Leakage Analysis
To compare the leakages of the masking and the baseline schemes, we proceed as
$,\displaystyle I(y_{1},$$\ldots,y_{M};d_{i}(x)+\mu(i),i\in[M]|x)$
$=$ $H(d_{1}(x)+\mu(1),\ldots,d_{M}(x)+\mu(M)|x)$
$-H(d_{i}(x)+\mu(i),i\in[M]|y_{1},\ldots,y_{M},x)$ (50)
$\leq$ $H(d_{1}(x),\mu(1),\ldots,d_{M}(x),\mu(M)|x)$
$-H(\mu(i),i\in[M]|y_{1},\ldots,y_{M},x)$ (51)
$=$ $H(d_{1}(x),\ldots,d_{M}(x)|x)$
$+H(\mu(1),\ldots,\mu(M)|x,d_{1}(x),\ldots,d_{M}(x))$
$-H(\mu(1),\ldots,\mu(M)|y_{1},\ldots,y_{M},x)$ (52)
$=$ $I(y_{1},\ldots,y_{M};d_{1}(x),\ldots,d_{M}(x)|x),$ (53)
where (53) is the leakage metric for the baseline scheme.

### v2

@@ III-C Mask-PCR
There exists a scheme that has a lower database leakage in terms of (5) compared to the baseline PCR with the communication cost of $2(d+M)$ symbols of $\mathbb{F}_{q}$ where $q>R^{2}d$ is prime.
@@
$d_{k}=\min_{i,j}|d_{i}(x_{k})-d_{j}(x_{k})|,\quad d_{\min}=\min_{k}d_{k}.$ (17)
Let $\mu$ be a random variable with support $\{0,\ldots,d_{\min}-1\}$. Now, a user who wishes to know the closest accepted sample to their rejected sample $x\in\mathcal{D}_{c}$ sends the query in (8) to the $n$th server. Upon receiving the queries, each server computes the answers as follows
$A_{n}^{[x]}(i)=||y_{i}-Q_{n}^{[x]}||^{2}+\mu(i)+\alpha_{n}Z^{\prime}(i),\quad i\in[M],$ (18)
where $\mu(i)$ has the same distribution as $\mu$ and $Z^{\prime}(i)$ is a uniform random variable in $\mathbb{F}_{q}$. As the user receives the answers from the servers, they are reprocessed as follows,
@@
and are used to decode the distances as follows,
@@
Upon getting the values of masked distances, $d_{i}(x)+\mu(i)$, $i\in[M]$, the user utilizes them to decide which is closest based on their numerical value, i.e., decode the index of the closest accepted sample. To show that the user can correctly decode with the masked distances, we need the following lemma, where its proof can be found in [29].
@@
Lemma 1 states that all the vectors in $\mathcal{D}_{c}$ preserve the relative distances among the accepted samples, i.e., the ordering of the distances among the points in $\mathcal{D}_{c}$ and accepted samples is maintained. In addition, note that the user does not know the exact value of $d_{\min}$ since they do not have any prior knowledge of the accepted samples.
Remark 3
We believe that the requirement $x\in\mathcal{D}_{c}$ is a bit strong, which is why we dropped this requirement in our first experiment (see Sec V-A) and we find $d_{\min}$ empirically instead. However, it is required for theoretical analysis to guarantee decodability.
Communication Cost
@@
Computation Complexity
The computational complexity at the user and the server is equal to that of baseline PCR.
Remark 4
It is important to note that (17) is evaluated prior to the scheme initiation and done once only till the dataset itself changes. This is a part of pre-processing for Mask-PCR and is not a part of the main scheme, which begins with the user sending their query. This is why it is not included in the calculation of the computation complexity.
Effect of Field Size in Masking
The field size can have a significant role in counterfactual retrieval with the masking approach. Recall that the minimum requirement for the field size is $R^{2}d$ as explained at the beginning of this section. Let $q_{1}$ be the field size satisfying $q_{1}>R^{2}d$. If another field size $q_{2}$ is chosen such that $q_{2}>q_{1}$, we can embed the samples $y_{1},\ldots,y_{M}$ in $\mathbb{F}_{q_{2}}$ using a relative distance preserving transform $T:\mathbb{F}_{q_{1}}\rightarrow\mathbb{F}_{q_{2}}$ such that the following condition is satisfied $\forall k,i,j$,
$|d_{i}(x_{k})-d_{j}(x_{k})|\leq$ $|\big{(}||x_{k}-T(y_{i})||^{2}-||x_{k}-T(y_{j})||^{2}\big{)}|.$ (21)
@@
This example demonstrates the masking approach and the field size expansion. Let the rejected samples, for example, be $\{[1,2]^{t},[2,1]^{t}\}$, and the set of accepted samples be $\{[20,0]^{t},[0,20]^{t}\}$. Let the field size be $q_{1}=809$. Thus, after simple calculations, we see that the range of the masking random variable is $\{0,\ldots,39\}$. Let the user choose $x=[1,2]^{t}$. Thus, the queries sent to the two servers are given by (8) and after the servers reply, the user reprocesses the received answers according to (19) to obtain
$A^{[x]}(1)=\begin{bmatrix}\hat{A}_{1}^{[x]}(1)\\ \hat{A}_{2}^{[x]}(1)\end{bmatrix}=\bm{M}_{2}\begin{bmatrix}365+\mu(1)\\ I(1)+Z^{\prime}(1)\end{bmatrix},$ (22)
@@
$A^{[x]}(2)=\begin{bmatrix}\hat{A}_{1}^{[x]}(2)\\ \hat{A}_{2}^{[x]}(2)\end{bmatrix}=\bm{M}_{2}\begin{bmatrix}325+\mu(2)\\ I(2)+Z^{\prime}(2)\end{bmatrix}.$ (23)
It is clear that for any choice of $\mu(1)$, and $\mu(2)$, the user would be able to know that $y_{2}$ is closest to $x$.
Now, let us choose a larger field size, for example, $q_{2}=40009$ and apply the simple transform $T:a\rightarrow 10a$. Then, with simple calculations, we see that the range of the masking random variable is $\{0,\ldots,399\}$ which is larger than the previous one and still maintains the relative distance.
Leakage Analysis
Similarly, on comparing the leakages of masking and baseline schemes, we show that the former is lower than the latter. The details of the proof are given in [29].
IV PCR WITH ACTIONABILITY
In this section, we assume that the user assigns different weights to their attributes based on their preference to change those attributes to attain a counterfactual. The more reluctant the user is to change a given attribute, the higher its weight is. In order to ensure that the user’s preference is also private, the weight vector should be kept private from the servers. Let $w\in[L]^{d}$ denote the user’s weight vector, where $L$ is a positive integer. Then, we have the following theorem.
@@ IV PCR WITH ACTIONABILITY
There exist schemes derived from Baseline PCR, Diff-PCR and Mask-PCR, that can retrieve the exact closest counterfactual with user’s actionability weights using $N=3$ servers while keeping the user’s sample and actionability weights hidden from each server. The communication costs of the respective schemes are:
  1. 1.
    Baseline PCR+: $6d+3M$ symbols of $\mathbb{F}_{q}$, $q>R^{2}Ld$
  2. 2.
    Diff PCR+: $6d+3(M-1)$ symbols of $\mathbb{F}_{q}$, $q>2R^{2}Ld$
  3. 3.
    Mask PCR+: $6d+3M$ symbols of $\mathbb{F}_{q}$, $q>R^{2}Ld$
where $q$ is prime.
To prove Theorem 4, we describe the extended schemes next.It is easy to verify that the PCR+ schemes guarantee actionability, at equal computational complexities as their PCR counterparts.
IV-A Baseline PCR+
In this case, the required minimum field size is $q\geq R^{2}Ld$. Now, the user sends the query tuple $Q_{n}^{[x,w]}$ to server $n$, where
$Q_{n}^{[x,w]}(1)=x+\alpha_{n}Z_{1},\quad Q_{n}^{[x,w]}(2)=w+\alpha_{n}Z_{2},$ (24)
@@ IV-A Baseline PCR+
$-2(y_{i}-x)^{t}\bm{W}Z_{1}+Z^{\prime}_{1}\Big{)}+\alpha_{n}^{2}\Big{(}Z_{1}^{t}\bm{W}Z_{1}$
$-2(y_{i}-x)^{t}\bm{Z}Z_{1}+Z^{\prime}_{2}\Big{)}+\alpha_{n}^{3}Z_{1}^{t}\bm{Z}Z_{1},$ (26)
@@
Upon receiving the answers from $N=3$ servers, the user applies the following decoding approach,
@@
$=\bm{M}_{3}\begin{bmatrix}d_{w}(y_{i},x),I_{1}(i),I_{2}(i)\end{bmatrix}^{t},$ (28)
@@
IV-B Diff-PCR+
@@
where $I_{1}^{\prime}(i)=(y_{i}-x)^{t}\bm{Z}(y_{i}-x)-(y_{i+1}-x)^{t}\bm{Z}(y_{i+1}-x)+2Z_{1}^{t}\bm{W}(y_{i}-y_{i+1})+Z^{\prime}_{1}(i)$ and $I^{\prime}_{2}(i)=2Z_{1}^{t}\bm{Z}(y_{i+1}-y_{i})+Z^{\prime}_{2}(i)$ are the interference terms. The answers $A_{n}^{[x,w]}(i)$ of the three servers can be written as,
$A^{[x,w]}(i)=\bm{M}_{3}\begin{bmatrix}d_{w}(y_{i},x)-d_{w}(y_{i+1},x)\\ I_{1}^{\prime}(i)\\ I_{2}^{\prime}(i)\end{bmatrix},$ (33)
where $A^{[x,w]}(i)=\begin{bmatrix}A_{1}^{[x,w]}(i),A_{2}^{[x,w]}(i),A_{3}^{[x,w]}(i)\end{bmatrix}^{t}$. Therefore, the user recovers the $M-1$ differences, while discarding the interference terms.
Finding the Closest $y_{i}$ to $x$: Now, with the $M-1$ differences, the user evaluates their counterfactual index $\theta^{*}$ following sequential comparisons similar to Algorithm 1, the only difference being the range in line 5 is replaced by $\mathcal{H}^{c}$.
@@
IV-C Mask-PCR+
The queries are the same as in the previous section. Upon receiving the queries, each server applies the following on each sample $y_{i}$ and sends to the user
@@ IV-C Mask-PCR+
where $\mu(i)$ is as defined in the previous section, $Z^{\prime}_{1}(i)$, and $Z^{\prime}_{2}(i)$ are uniform random variables and $I_{1}(i)$ and $I_{2}(i)$ are as given in (IV-A). Using the answers of the $N=3$ servers and applying the decoding approach given in Section IV-A, the user obtains the masked weighted distance corresponding to $y_{i}$ as follows:
$\hat{A}^{[x,w]}(i)=\bm{M}_{3}\begin{bmatrix}d_{w}(y_{i},x)+\mu(i),I_{1}(i),I_{2}(i)\end{bmatrix}^{t}.$ (36)
To show that there is no need to change the range of the masking random variable $\mu$ with any weighting, we provide the following lemma, which is proven in [29].
@@
V EXPERIMENTS
V-A Accuracy-Quantization Trade-Off
As detailed in Section III-A, the proposed schemes operate on finite fields with size $q$. However, a real-world application might require some of the features to be real-valued. To circumvent this technicality, we may quantize each feature of the counterfactual instances as well as the user sample before applying the schemes to align with our system model. However, the quantization of each feature to integers in $[0:R]$ will lead to a loss of some information that might result in an error in finding the closest counterfactual. The error can be made arbitrarily small by increasing the number of quantization levels, as we demonstrate in the following experiments. The trade-off is that a larger finite field $\mathbb{F}_{q}$ will be required for the scheme to work properly; e.g., for two implementations of Mask-PCR with numbers of quantization levels $R_{1},R_{2}$ such that $R_{1}>R_{2}$, the minimum field sizes required are $q_{1}\geq q_{2}$ because $q_{1}>R_{1}^{2}d$ and $q_{2}>R^{2}_{2}d$.
Moreover, in the case of Mask-PCR, we may lift the restriction $x\in\mathcal{D}_{c}$ off by empirically deciding the parameter $d_{\min}$ to be used such that the scheme works for most of the points in a sample set of queries. As pointed out in Section III-C, a larger field size $q$ is needed to accommodate a larger value for $d_{\min}$.
We carry out our experiment using the Wine Quality Dataset [30] to observe the accuracy-quantization trade-offs mentioned above. The dataset contains 4898 instances, each with 11 real-valued features. The original target variable “Quality” is categorical, taking integer values from 0 to 10. We convert the target into a binary variable by defining $\tau=\mathds{1}[``\text{Quality}"\geq 5]$. This gives us 3788 instances with $\tau=1$ and 183 instances with $\tau=0$ (which are the potential queries).
We consider all the instances with $\tau=1$ to be the accepted instances (potential counterfactual instances). Let this set be $\mathcal{S}_{1}$. The set $\mathcal{S}_{0}$ consists of all the instances with $\tau=0$ which are the potential queries. In each round, we pick a subset $\mathcal{D}$, with $|\mathcal{D}|=M$, of $\mathcal{S}_{1}$ as the database and a subset $\mathcal{S}_{x}$ of $\mathcal{S}_{0}$ as the queries. Then, we apply the retrieval schemes with varying parameters to obtain the nearest neighbor counterfactuals. The accuracy of the retrieved counterfactuals is computed as follows: Denote the counterfactual of the query $x$ retrieved using the scheme by $\hat{y}(x)$ and the actual (i.e., the exact nearest neighbor) counterfactual by $y(x)$. Note that the scheme provides the index of the counterfactual, hence, we can retrieve the exact unquantized counterfactual irrespective of the quantization applied. The accuracy of the scheme is
@@ V-A Accuracy-Quantization Trade-Off
Fig. 2, shows how the accuracy improves with an increasing number of quantization levels. Further, for a given number of quantization levels, the accuracy degrades as $d_{\min}$ increases.
@@
V-B Database Leakage Results
V-B1 Synthetic Dataset
To observe the success of Diff-PCR and Mask-PCR in mitigating the leakage, we compute the exact leakage values over a synthetic dataset. We refer to [29] for leakage computation. We consider $R=3$ and $d=3$ with a database size of $M=3$. We assume the queries are equi-probable over the $[0:R]^{d}$ space. We further assume that given the query $x$, $y_{1},\dots,y_{M}$ are equi-probable over $[0:R]^{d}\backslash\{x\}$. Table I lists the computed leakage values under these assumptions.
Scheme Leakage (to base $q=757$)
Baseline PCR 1.143
Diff-PCR 0.940
Mask-PCR ($d_{\text{min}}=2$) 0.911
Mask-PCR ($d_{\text{min}}=3$) 0.777
TABLE I: Leakage results – synthetic data.
V-B2 COMPAS dataset
Correctional Offender Management Profiling for Alternative Sanctions (COMPAS) dataset [31] includes 6172 instances and 20 features (after one-hot encoding the categorical features) about individuals, which are used to predict the likelihood of reoffending of a convicted criminal. The target variable $\tau$ is “is_recid”. Class-wise counts are 3182 and 2990 for $\tau=0$ and $\tau=1$, respectively. We normalize all the features to the interval $[0,1]$ and quantize to $R+1$ levels with $R=10$ during pre-processing, leaving 560 unique instances with $\tau=0$ and 717 unique instances with $\tau=1$. As our schemes are independent of the classifier, we use these labels as the classifications of the instances. In order to compute the leakage values, we assume the following:
  1. 1.
    Queries are equi-probable over the $\tau=0$ (rejected) instances.
  2. 2.
    Since computing the histogram over all possible $M$-tuples (with $M=5$ in this particular experiment) is too computationally intensive, we sample $10^{5}$ of the possible permutations and evaluate the leakage over those. Accordingly, given the query $x$, the $M$-tuples $y=(y_{1},\dots,y_{M})$ are assumed to be equi-probable over a set of $10^{5}$ $M$-tuples that can be generated using the $\tau=1$ (accepted) instances.
Table II presents the results. Similar to the values obtained with the synthetic dataset, we observe that Diff-PCR and Mask-PCR schemes reduce leakage compared to the baseline scheme.
Scheme Leakage (to base $q=2003$)
Baseline PCR 1.51437
Diff-PCR 1.51436
Mask-PCR ($d_{\text{min}}=2$) 1.51427
TABLE II: Leakage results – COMPAS dataset.